# The Architecture and Evaluation of Bayesian Neural Networks

### v1

@@ 4 Discussion
The message of an optimist’s conclusion could question the common belief that the mean-field variational approximations are generally overly restrictive and do not capture the true posterior and the uncertainty well. Even given the increased available computing power and assuming that HMC can be taken as a gold standard (which as we have seen is not the case), computational costs of the sampling algorithms suggest that it may not be feasible for most of the modern neural networks and datasets. Indeed, in a variety of the experiments considered in Sections 2.3, 2.4, 2.5 and 3.3 mfVI overall provided better uncertainty quantification than HMC, and in out-of-distribution setting the empirical coverage of the latter was close to zero. We note that for single-layer neural networks HMC outperformed mfVI only in terms of accuracy. At the same time, for deeper networks and in out-of-distribution scenarios, the accuracy of mfVI was often comparable to HMC. Further, in Section 2.4 we confirmed that even for slightly deeper networks the time needed to perform HMC becomes a burden which makes variational inference a very attractive alternative to sampling. Nevertheless, in Section 2.3 we observed that the restrictions imposed by the factorized families can obstruct models from effectively learning from the data. In real-life scenarios where one is required to evaluate the future predictive performance of the model before applying it to the unseen data, the estimate of the expected log pointwise predictive density can serve as a reliable diagnostic and thus, PSIS-LOO estimates can be beneficial for model assessment and combination. In Section 3.3, stacking as well as, ensembles of BNNs were shown to be a possible solution when dealing with multimodal posteriors, helping to both improve accuracy and uncertainty quantification even in the extreme OOD scenario. We find that stacked or ensembled variational approximations are competitive to HMC at a much-reduced cost.
@@
This work highlights the model’s sensitivity to architectural choices, namely, width, depth and activation function. Along the same line, future work could study the performance of various priors, including sparsity-inducing priors which are known to reduce computational costs whilst improving the accuracy and calibration [Blundell et al., 2015, Polson and Ročková, 2018]. Further, an important avenue for research is to consider the so-called structured variational inference with less restrictive variational families, and more generally, study the trade-off between the expressiveness of the variational family and scalability. Finally, given the multimodal nature of distributions arising in Bayesian neural networks, a promising avenue for research is to continue improving model combination techniques, possibly along the lines of an adaptive variational Bayes framework [Ohn and Lin, 2024] or hierarchical stacking and combining the models pointwise [Yao et al., 2022].
@@ Appendix A Metrics and Practicalities
The results are obtained when the number of iterations of mfVI is set $10^{4}$, $10^{4}$, $5\times 10^{4}$, $6\times 10^{4}$ to train models with, respectively, $20,200,1000,2000$ hidden units in a layer. In models trained with HMC, the number of samples used for warmup was set to $10^{3}$, the samples used for posterior is $10^{3}$ in models with $20$ hidden units, and $2\times 10^{3}$ in models with $200,1000,2000$ hidden units in a layer. In experiments the number of iterations of mfVI was set to $L\times 10^{4}$; the HMC had the number of warmup samples fixed to $10^{3}$ and the number of samples was $\min(4\times 10^{3},L\times 10^{3})$.
@@ B.1 Empirical model assessment
Recall, the OOD scenario, where we have specifically created an extreme case when the training data $\bm{\mathcal{D}}_{c}$ and the new data $\tilde{\bm{\mathcal{D}}}_{c}$ were covering disjoint regions of the true function. To assess the expected out-of-sample predictive performance, we could begin with posterior predictive checks (PPC), which compare the true $\bm{\mathcal{D}}_{c}$ to datasets simulated from the posterior predictive distribution Gelman et al. [2020]. Figure 6(a) provides the PPC based on the kernel density estimates of the observed $\bm{y}$ and Figure 6(b) compares posterior predictive to the observed $\bm{y}$ for all of the models with $D_{1}=2000$; both figures evidently classify the mean-field VI with sigmoid activation as inappropriate. However, it is not apparent that HMCS2000 is significantly inferior to HMCR2000. We compare the RMSE computed in the OOD settings to $\widehat{\text{elpd}}_{\text{loo}}$ estimates. Since the idea of estimating the expected log predictive density is in evaluating future predictive performance, we expect the higher $\widehat{\text{elpd}}_{\text{loo}}$ to correspond to a better model and lower RMSE. Indeed, Figure 6(c) is more informative in this sense than PPC diagnostics, and we observe the inverse dependence between $\widehat{\text{elpd}}_{\text{loo}}$ and RMSE. Sampling and approximation techniques result in different scales of $\widehat{\text{elpd}}_{\text{loo}}$ estimates (in general, $\widehat{\text{elpd}}_{\text{loo}}$ is lower for VI, especially, in wide networks) and thus, we compare the models trained with different algorithms for better visualization.
@@
Figure 6: Estimating the out-of-distribution performance before seeing the new data: testing the (a), (b) PPC and (c) $\widehat{\text{elpd}}_{\text{loo}}$. The mfVIR2000 is confirmed to be unreliable in all methods. The PPC of the HMCS2000 does not provide enough information to judge its performance in the OOD settings, while the $\widehat{\text{elpd}}_{\text{loo}}$ does.
@@ B.2 Correspondence Between WAIC and RMSE
With $\widehat{\text{lpd}}$ at hand, a superior successor of the Akaike Information criterion (AIC) Akaike [1998] and the Deviance information criterion (DIC) Spiegelhalter et al. [2002] called the Watanabe-Akaike information criterion (WAIC)  [^2] Also called Widely Applicable information criterion Watanabe [2010] can be obtained. To mitigate the bias between the log pointwise density and the expected utility, WAIC subtracts the simulation-estimated effective number of parameters:
@@
Here, $\operatorname{Var}^{S}$ is the sample variance and the estimated effective number of parameters $\widehat{p}_{\text{WAIC}}$ can be seen as a measure of model complexity. Asymptotically, WAIC is equivalent to the Bayesian leave-one-out cross-validation (LOO-CV) estimate of the expected utility Watanabe [2010]. Even though cross-validation is a natural framework for accessing the model’s predictive performance, the WAIC was for a long time preferred over the LOO-CV due to the computational challenges arising from multiple model runs Gelman et al. [2014]. Moreover, whilst both the PSIS-LOO and WAIC estimates give nearly unbiased estimates of the predictive ability of the model, $\widehat{\text{elpd}}_{\text{loo}}$ was shown to be more robust than $\widehat{\text{elpd}}_{\text{WAIC}}$; in the presence of limited sample size and weak priors, WAIC can severely underestimate $\hat{p}_{\text{WAIC}}$ and often has a larger bias towards the log predictive density Vehtari et al. [2016], Gelman et al. [2014].
That said, computing $\widehat{\text{elpd}}_{\text{loo}}$ involved Pareto smoothed importance sampling, and in some of the models, the estimated shape parameter of the generalized Pareto distribution gave a warning about the reliability of the LOO estimate. Here we do an extra step and check if the WAIC and LOO estimates of the elpd agree. Figure 7 illustrates the reverse dependency between the RMSE and $\widehat{\text{elpd}}_{\text{WAIC}}$ and is largely identical to the figure illustrating the dependency between the RMSE and $\widehat{\text{elpd}}_{\text{loo}}$ (in terms of the location of coordinates but not the error bars). Therefore, we can conclude that LOO estimates of the elphd can be seen as relatively reliable.
Appendix C Supplementary to Ensembles and Averages
@@
C.1 Milder OOD Scenario
Given the nature of the test data we use, it is reasonable to assume that not only the predictions but also the $\widehat{\text{elpd}}_{\text{loo}}$ estimates are tentative. Thus, in order to explore different methodologies from different angles we create a simpler regression problem in which test data comes from a slightly broader region. Using the classification of Farquhar and Gal [2022], this new scenario could be named as an OOD task with ’related-distributions’. We define a similar synthetic dataset with one-dimensional input and output, to which we add some small noise:
$\bm{x}$ $\sim\text{Unif}([0,1])$
$\bm{y}$ $=\sin(10\bm{x})\bm{x}^{2}+\bm{\epsilon}$
$\bm{\epsilon}$ $\sim 0.05\text{N}(0,1)$
As before, the input is scaled, unlike the output. The data for training $\bm{\mathcal{D}}_{r}$ and the testing $\tilde{\bm{\mathcal{D}}}_{r}$ consist of $N=450$ and $\tilde{N}=50$ observations, respectively, where $\tilde{\bm{\mathcal{D}}}_{r}$ comes from the broader than $\bm{\mathcal{D}}_{r}$ region, i.e. $(\min_{n=1\ldots N}(x_{n}),\max_{n=1\ldots N}(x_{n}))\subset(\min_{n=1\ldots\tilde{N}}(\tilde{x}_{n}),\max_{n=1\ldots\tilde{N}}(\tilde{x}_{n}))$. Having 10 posterior predictive distributions of mfVIR20, we compare ensembling, pseudo-BMA and stacking on Figure 8. Whilst the total uncertainty estimates of pseudo-BMA are, somewhat, adequate, the model uncertainty is underestimated. Stacking provides slightly better accuracy and empirical coverage than the deep ensembles, with both techniques leading to improved prediction performance and uncertainty quantification.
@@ C.1 Milder OOD Scenario
Figure 8: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 in the ’related-distributions’ task. The pseudo-BMA is again worse than the other methodologies. Stacking is slightly better than DE.
C.2 HMC and Averaging Techniques
We wish to recreate the experiment for the HMCR20 model (instead of the mfVIR20 mode). We choose 10 random initialization points, obtain 10 posterior predictive distributions and compute estimated expected log pointwise predictive densities. We then construct ensemble, pseudo-BMA and stacking approximations, for the ’complement-distributions’ task the results are illustrated by Figure 9 and the predictions of the ’related-distributions’ task are shown on Figure 10. In contrast to mfVIR20, this time we do not observe a clear difference between the pseudo-BMA and stacking and ensembling methodologies. Moreover, all the approaches require considerable computational resources (for 10 random runs) but do not provide a considerable improvement in RMSE and empirical coverage compared to a single random run of the model. We conclude that in this particular example, ensembling, stacking and pseudo-BMA do not help explore different modes of the posterior and so cannot be recommended when dealing with HMC.
Figure 9: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to HMCR20 in the ’complement-distributions’ task are very similar.
Figure 10: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to HMCR20 in the ’related-distributions’ task are (again) strikingly similar.
C.3 Deeper Networks
Based on the 10 posterior predictive distributions obtained starting from 10 different random initialization points, we construct ensemble, pseudo-BMA and stacking approximations for mfVIR and mfVIS models with $L=6$ hidden layers. The results are consistent with the observation made in the main body of the work; in the ’complement-distributions’ task (Figure 12) pseudo-BMA is confirmed to be inferior to stacking and deep ensembles of BNNs. In the related distribution task (Figure 11), in terms of both accuracy and uncertainty quantification, stacking is preferable over deep ensembles and pseudo-BMA with the latter performing better than ensembles (unlike in the one-layer case).
@@ C.3 Deeper Networks
Figure 11: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with $L=6$ in the ’complement-distributions’ task.
Figure 12: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with $L=6$ in the ’related-distributions’ task.

### v2

@@ 4 Discussion
The message of an optimist’s conclusion could question the common belief that the mean-field variational approximations are generally overly restrictive and do not capture the true posterior and the uncertainty well. Even with increases in computing power, the computational costs of sampling algorithms suggest that it may not be feasible for most modern neural networks and datasets. Moreover, although HMC is often considered as a gold standard, we have seen this may not be the case for BNNs due to complexity and multimodality of the posterior. Indeed, in a variety of experiments considered in Sections 2.3, 2.4, 2.5 and 3.3 mfVI overall provided better uncertainty quantification than HMC, and in out-of-distribution settings, the empirical coverage of the latter was close to zero. We note that for single-layer neural networks, HMC outperformed mfVI only in terms of accuracy. At the same time, for deeper networks and in out-of-distribution scenarios, the accuracy of mfVI was often comparable to HMC. Further, in Section 2.4 we confirmed that even for slightly deeper networks the time needed for HMC becomes a burden, which makes variational inference a very attractive alternative to sampling. Nevertheless, in Section 2.3 we observed that the restrictions imposed by the factorized families can obstruct models from effectively learning from the data. In real-life scenarios where one is required to evaluate the future predictive performance of the model before applying it to the unseen data, the estimate of the expected log pointwise predictive density can serve as a reliable diagnostic and thus, PSIS-LOO estimates can be beneficial for model assessment and combination. In Sections 3.3 and 3.4, stacking and ensembles of BNNs were shown to be a possible solution when dealing with multimodal posteriors, helping to both improve accuracy and uncertainty quantification even in the extreme OOD scenario. We find that stacked or ensembled variational approximations are competitive to HMC at a much-reduced cost. Not only do the performance of individual models depend on architectural choices, but as observed in Section 3.4, the model averaging techniques are themselves influenced by these modelling choices. Finally, we note that overall in our experiments, there was no considerable and systematic difference in the performance between the BNNs with Gaussian and Student-t priors.
@@
This work highlights the model’s sensitivity to architectural choices, namely, width, depth and activation function. Future work could study the performance of various more elaborate than Gaussian or Student-t choices of priors placed on the weights, including sparsity-inducing priors which have been shown to improve the accuracy and calibration [Blundell et al., 2015, Polson and Ročková, 2018]. Further, an important avenue for research is to consider the so-called structured variational inference with less restrictive variational families, and more generally, study the trade-off between the expressiveness of the variational family and scalability. Finally, given the multimodal nature of distributions arising in Bayesian neural networks, a promising avenue for research is to continue improving model combination techniques. This includes developing a better understanding of the number of models required for optimal performance with existing ensembling methods, as well as exploring more advanced approaches such as adaptive variational Bayes frameworks [Ohn and Lin, 2024] or hierarchical stacking and pointwise model combination [Yao et al., 2022].
@@ Appendix A Metrics and Practicalities
The results are obtained when the number of iterations of mfVI is set $10^{4}$, $10^{4}$, $5\times 10^{4}$, $6\times 10^{4}$ to train models with, respectively, $20,200,1000,2000$ hidden units in a layer. In models trained with HMC, the number of samples used for warmup was set to $10^{3}$, the samples used for posterior is $10^{3}$ in models with $20$ hidden units, and $2\times 10^{3}$ in models with $200,1000,2000$ hidden units in a layer. Across all the experiments with simulated data, the learning rate of the optimizer is set to $35e-$ 3. In deeper networks, the number of iterations of mfVI was set to $L\times 10^{4}$; the HMC had the number of warmup samples fixed to $10^{3}$ and the number of samples was $\min(4\times 10^{3},L\times 10^{3})$.
When evaluating BNN’s performance on the NASA’s rocket booster simulation, we only considered mfVI, and set the number of iterations to $10^{4}$ and learning rate to $25e-$ 2, with one exception of the OOD scenario and 200 hidden units, when we trained the model for $2\times 10^{4}$ iterations.
@@ B.1 Empirical model assessment
Recall, the OOD scenario, where we have specifically created an extreme case when the training data $\bm{\mathcal{D}}_{c}$ and the new data $\tilde{\bm{\mathcal{D}}}_{c}$ were covering disjoint regions of the true function. To asses the expected out-of-sample predictive performance, we could begin with posterior predictive checks (PPC), which compare the true $\bm{\mathcal{D}}_{c}$ to datasets simulated from the posterior predictive distribution [Gelman et al., 2020]. Figure 8(a) and Figure 9(a) provide the PPC based on the kernel density estimates of the observed $\bm{y}$, and Figure 8(b) and Figure 9(b) compare posterior predictive to the observed $\bm{y}$ for all of the models with $D_{1}=2000$; both figures evidently classify the mean-field VI with sigmoid activation as inappropriate. However, it is not apparent that HMCS2000 is significantly inferior to HMCR2000. We compare the RMSE computed in the OOD settings to $\widehat{\text{elpd}}_{\text{loo}}$ estimates. Since the idea of estimating the expected log predictive density is in evaluating future predictive performance, we expect the higher $\widehat{\text{elpd}}_{\text{loo}}$ to correspond to a better model and lower RMSE. Indeed, Figure 8(c) and Figure 9(c) are more informative in this sense than PPC diagnostics, and we observe the inverse dependence between $\widehat{\text{elpd}}_{\text{loo}}$ and RMSE. Sampling and approximation techniques result in different scales of $\widehat{\text{elpd}}_{\text{loo}}$ estimates (in general, $\widehat{\text{elpd}}_{\text{loo}}$ is lower for VI, especially, in wide networks) and thus, we compare the models trained with different algorithms for better visualization.
@@
Figure 8: Estimating the out-of-distribution performance of models with Gaussian priors before seeing the new data: testing the (a), (b) PPC and (c) $\widehat{\text{elpd}}_{\text{loo}}$. The mfVIR2000 is confirmed to be unreliable in all methods. The PPC of the HMCS2000 does not provide enough information to judge its performance in the OOD settings, while the $\widehat{\text{elpd}}_{\text{loo}}$ does.
Figure 9: Estimating the out-of-distribution performance of models with Student-t priors before seeing the new data: testing the (a), (b) PPC and (c) $\widehat{\text{elpd}}_{\text{loo}}$. The mfVIR2000 is confirmed to be unreliable in all methods. The PPC of the HMCS2000 does not provide enough information to judge its performance in the OOD settings, while the $\widehat{\text{elpd}}_{\text{loo}}$ does.
@@ B.2 Correspondence Between WAIC and RMSE
With $\widehat{\text{lpd}}$ at hand, a superior successor of the Akaike Information criterion (AIC) [Akaike, 1998] and the Deviance information criterion (DIC) [Spiegelhalter et al., 2002] called the Watanabe-Akaike information criterion (WAIC)  [^2] Also called Widely Applicable information criterion [Watanabe, 2010] can be obtained. To mitigate the bias between the log pointwise density and the expected utility, WAIC subtracts the simulation-estimated effective number of parameters:
@@
Here, $\operatorname{Var}^{S}$ is the sample variance and the estimated effective number of parameters $\widehat{p}_{\text{WAIC}}$ can be seen as a measure of model complexity. Asymptotically, WAIC is equivalent to the Bayesian leave-one-out cross-validation (LOO-CV) estimate of the expected utility [Watanabe, 2010]. Even though cross-validation is a natural framework for accessing the model’s predictive performance, the WAIC was for a long time preferred over the LOO-CV due to the computational challenges arising from multiple model runs [Gelman et al., 2014]. Moreover, whilst both the PSIS-LOO and WAIC estimates give nearly unbiased estimates of the predictive ability of the model, $\widehat{\text{elpd}}_{\text{loo}}$ was shown to be more robust than $\widehat{\text{elpd}}_{\text{WAIC}}$; in the presence of limited sample size and weak priors, WAIC can severely underestimate $\hat{p}_{\text{WAIC}}$ and often has a larger bias towards the log predictive density [Vehtari et al., 2016, Gelman et al., 2014].
That said, computing $\widehat{\text{elpd}}_{\text{loo}}$ involved Pareto smoothed importance sampling, and in some of the models, the estimated shape parameter of the generalized Pareto distribution gave a warning about the reliability of the LOO estimate. Here, we do an extra step and check if the WAIC and LOO estimates of the elpd agree in the case of models with Gaussian priors. Figure 10 illustrates the reverse dependency between the RMSE and $\widehat{\text{elpd}}_{\text{WAIC}}$ and is largely identical to the figure illustrating the dependency between the RMSE and $\widehat{\text{elpd}}_{\text{loo}}$ (in terms of the location of coordinates but not the error bars). Therefore, we can conclude that LOO estimates of the elphd can be seen as relatively reliable.
Appendix C Supplementary for Student-t Priors
Figure 11: The predictions and uncertainty estimates obtained by each model for $D_{1}=2000$ and Student-t priors.
In the main body of the work, we first studied the performances of mfVIR, mfVIS, HMCR and HMCS with 1 hidden layer and either Gaussian or Student-t priors as the width increases. The predictions of all four models with Student-t priors when $D_{1}=2000$ are provided on the Figure 11. Similar to models with Gaussian priors, the performance of the mfVIS dips with the increase in the dimension of the hidden layer; for wider networks, its posterior predictive distribution fails to capture the data, and degenerates to the prior. The HMCR2000 model exhibits a better performance than the HMCS2000 both in terms of accuracy and uncertainty quantification.
Consider BNNs with Student-t priors with the number of layers $L$ varying from 1 to 6 and a fixed number of hidden units in each layer $D_{h}=20$, then Figure 12 illustrates the predictions of all four models with $L=6$
Figure 12: The predictions and uncertainty estimates obtained by each model for $L=6$ and Student-t priors.
Further, Figure 13 compares non-OOD and OOD predictions obtained by the BNNs with ReLU activation, Student-t priors and $D_{1}=200$.
Figure 13: Within-the sample and out-of-sample predictions and uncertainty estimates of BNNs with $D_{1}=200$, ReLU activation and Student-t priors.
Appendix D Supplementary to Ensembles and Averages
@@
D.1 Models with Student-t priors
@@
Figure 14: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with Student-t priors in the ’complement-distributions’ (top) and ’related-distributions’ (bottom) OOD tasks. The pseudo-BMA is worse than DE and stacking, which are very similar to each other.
Consider the mfVIR20 model with Student-t priors, the ’complement-distributions’ and the ’related-distributions’ OOD tasks. In each task, we obtain 10 posterior predictive distributions and construct ensemble, pseudo-BMA and stacking approximations, the results are illustrated Figure 14.
D.2 Deeper Networks
@@
Figure 15: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with $L=6$ in the ’complement-distributions’ (top) and the ’related-distributions’ (bottom) task.
Based on the 10 posterior predictive distributions obtained starting from 10 different random initialization points, we construct ensemble, pseudo-BMA and stacking approximations for mfVIR and mfVIS models with Gaussian priors and $L=6$ hidden layers. The results are consistent with the observation made in the main body of the work; in the ’complement-distributions’ task (Figure 15) pseudo-BMA is confirmed to be inferior to stacking and deep ensembles of BNNs. In the related distribution task (Figure 15), in terms of both accuracy and uncertainty quantification, stacking is preferable over deep ensembles and pseudo-BMA with the latter performing better than ensembles (unlike in the one-layer case).
D.3 Supplementary to the Rocket Booster Simulation
Figure 16: Slices of predicted lift obtained by a single model and by ensembling, stacking and pseudo-BMA when applied to mfVIR100 and mfVIR200 with Gaussian priors in the non-OOD (top figures) and ’complement-distributions’ OOD (bottom figures) tasks. Stacking outperforms other methodologies.
When considering the Langley Glide-Back Booster simulator data [Rogers et al., 2003], we evaluated the performance of the mfVIR for various widths and in both non-OOD and ’complement-distributions’ tasks. Here we provide additional figures for models with 100 and 200 hidden units illustrating the slices of predicting lift, with Mach on the x-axis and angles $\alpha$ and $\beta$ fixed to 25 and 4, respectively.